### v1

@@ 1 Introduction
In the protein domain, which still enjoys a substantial advantage in number of solved structures over RNA, deep learning approaches have made significant advances in protein property prediction [1] and design [2]. Along with larger datasets, the methods developed for proteins were propelled by the development of sophisticated representations of structures which exposed them to geometric deep learning methods such as graph neural networks [3], and large-scale self-supervised training [4].
RNAs are distinct from proteins and have their own specificity. In particular, they fold hierarchically. A secondary structure made of Watson-Crick and Wobble base pairs is rapidly assembled to serve as a scaffold for 3D structures [5]. During that process, non-canonical base interactions [6] are formed to stabilize the tertiary architecture. This information, along with others that will be discussed below, provides rich graphical models that can be used to develop customized solutions for analyzing and designing RNA structures.
Building on these customized solution and increasingly available RNA structural data, opportunities to uncover structure-function relationships and validate design algorithms through data-driven approaches for RNA come within reach [7]. With rnaglib [8], we take a step towards these advances by exposing a powerful graph encoding of RNA 3D structures, providing self-supervised learning training regimes, and exposing convenient dataset construction utilities to lower the barrier to entry.
@@ 1.2 Representing RNA 3D structures
In order to work with deep learning methods, we need a biophysical model compatible with state of the art data-driven methods. Specifically for RNA, the set of basepair interactions naturally leads to a graph representation where each node is a base, and an edge is a basepair. For ligand binding, it is known that detailed geometric orientations of binding sites is crucial and the typical secondary structure encoding can often lose this signal [7]. For this reason, a promising approach is to use the Leontis-Westhof classification which is made available through software packages such as x3dna-dssr [9] and FR3D [10]. See Fig. 1 top panel for an overview of the representation learning workflow.
@@
Having built an expressive encoding of the RNA 3D structure, that is here denoted as RNA 2.5D structure, we can leverage encoder-decoder representation learning methods [11] to predict functional attributes of interest (e.g. small-molecule binding, chemical modifications, protein binding, etc.) as shown in Figure 1. Here, we rely on an $\textrm{ENC}_{\phi}$ model parametrized by $\phi$, typically a neural network $\rightarrow\mathbb{R}^{d}$ which map nodes in a graph to real-valued embedding vectors $\mathbf{z}_{u}$ for each node $u\in G$. At this stage, the encoder can already be trained trained through self-supervised paradigm such as those which enforce distances in the embedding space to reflect those of a domain-specific similarity function on the graph space. For more details on this paradigm we point the reader to [12] and note that advanced functionalities of rnaglib provide support for self supervised learning but are not covered in this chapter. Adding a trainable output layer ($\textrm{DEC}_{\psi}$) to the RNA embedding models which is tuned for a given dataset allows pre-trained models to perform downstream property prediction tasks.
@@ 1.3 Objectives
The objective of this tutorial is to familiarize RNA and machine learning practitioners with the functionalities of rnaglib, which support RNA 3D dataset access and construction as well as machine learning model training and evaluation.
@@ 3.1.1 Obtaining annotated structures
All data is downloaded by default to ~/.rnaglib unless specified otherwise with the --download_dir flag. By specifying --redundancy [nr, all], you select whether to download a complete copy of all annotated 3D structures available at the time of building in the PDB or, with --redundancy nr, a representative set built by the BGSU RNA 3D Hub [10]. The --version specifies which release of the datasets to use with higher version numbers reflecting more recent builds.
@@ 3.2 Inspecting the RNA 3D object
Each RNA is stored as a dictionary with each key providing some basic information about the molecule. Upon creation, the dictionary contains the ID, and a networkx graph object of its interactions, and a path to the source data file.
@@ 4.2 Training Loop
4.3 Visualizing Results
We can return to the dataset to draw one of the graphs using the drawing package (See Fig. LABEL:fig:draw for a sample output).
@@ 4.3 Visualizing Results
Note that the layout is not optimized for RNA and is intended simply provided for checks and debugging. A list of node colors can be passed with the node_colors argument which can be used to visualize the model outputs on the RNA. By default, nodes are colored according to nucleotide identity.
@@ 5 Conclusion
In this chapter we demonstrated the use of rnaglib for convenient access to annotated RNA 3D structures and its use in a property prediction deep learning pipeline. We hope this illustration will lower the barrier to entry into the emerging field of geometric deep learning on RNA 3D structures and help inform novel RNA designs and uncover new structure-function relationships from 3D structure datasets. Finally, we point the reader again to the package documentation for a complete description of functionalities, including modules for self-supervised training and evaluation, and offer an invitation to contribute new features to the tool.

### v2

@@ 1 Introduction
In the protein domain, which still enjoys a substantial advantage in number of solved structures over RNA, deep learning approaches have made significant advances in protein property prediction [deepfri] and design [baker]. Along with larger datasets, the methods developed for proteins were propelled by the development of sophisticated representations of structures which exposed them to geometric deep learning methods such as graph neural networks [graphgen], and large-scale self-supervised training [esm].
RNAs are distinct from proteins and have their own specificity. In particular, they fold hierarchically. A secondary structure made of Watson-Crick and Wobble base pairs is rapidly assembled to serve as a scaffold for 3D structures [how]. During that process, non-canonical base interactions [lw] are formed to stabilize the tertiary architecture. This information, along with others that will be discussed below, provides rich graphical models that can be used to develop customized solutions for analyzing and designing RNA structures.
Building on these customized solution and increasingly available RNA structural data, opportunities to uncover structure-function relationships and validate design algorithms through data-driven approaches for RNA come within reach [rnamigos]. With rnaglib [rnaglib], we take a step towards these advances by providing a powerful graph encoding of RNA 3D structures, self-supervised learning training regimes, and dataset construction utilities to lower the barrier to entry.
@@ 1.2 Representing RNA 3D structures
In order to work with deep learning methods, we need a biophysical model compatible with state of the art data-driven methods. Specifically for RNA, the set of basepair interactions naturally leads to a graph representation where each node is a base, and an edge is a basepair. For ligand binding, it is known that detailed geometric orientations of binding sites is crucial and the typical secondary structure encoding can often lose this signal [rnamigos]. For this reason, a promising approach is to use the Leontis-Westhof classification which is made available through software packages such as x3dna-dssr [dssr] and FR3D [fr3d]. See Fig. 1 top panel for an overview of the representation learning workflow.
@@
Having built an expressive encoding of the RNA 3D structure, that is here denoted as RNA 2.5D structure, we can leverage encoder-decoder representation learning methods [ham] to predict functional attributes of interest (e.g. small-molecule binding, chemical modifications, protein binding, etc.) as shown in Figure 1. Here, we rely on an $\textrm{ENC}_{\phi}$ model parametrized by $\phi$, typically a neural network $\rightarrow\mathbb{R}^{d}$ which map nodes in a graph to real-valued embedding vectors $\mathbf{z}_{u}$ for each node $u\in G$. At this stage, the encoder can already be trained through self-supervised paradigm such as those which enforce distances in the embedding space to reflect those of a domain-specific similarity function on the graph space. For more details on this paradigm we point the reader to [vernal] and note that advanced functionalities of rnaglib provide support for self supervised learning but are not covered in this chapter. Adding a trainable output layer ($\textrm{DEC}_{\psi}$) to the RNA embedding models which is tuned for a given dataset allows pre-trained models to perform downstream property prediction tasks.
@@ 1.3 Objectives
The objective of this tutorial is to familiarize RNA informatics and machine learning practitioners with the functionalities of rnaglib, which support RNA 3D dataset access and construction as well as machine learning model training and evaluation.
@@ 3.1.1 Obtaining annotated structures
All data is downloaded by default to ~/.rnaglib unless specified otherwise with the --download_dir flag. By specifying --redundancy [nr, all], you select whether to download a complete copy of all annotated 3D structures available at the time of building in the PDB or, with --redundancy nr, a representative set built by the BGSU RNA 3D Hub [fr3d]. The --version specifies which release of the datasets to use with higher version numbers reflecting more recent builds.
@@ 3.2 Inspecting the RNA 3D object
Each RNA is stored as a dictionary with each key mapping to some basic information about the molecule. Upon creation, the dictionary contains the ID, and a networkx graph object of its interactions, and a path to the source data file.
@@ 4.2 Training Loop
4.3 Visualizing Structures
We can return to the dataset to draw one of the graphs using the drawing package (See Fig. 2 for a sample output).
@@ 4.3 Visualizing Structures
Note that the layout is not heavily optimized for RNA and is intended simply provided for checks and debugging. A list of node colors can be passed with the node_colors argument which can be used to visualize the model outputs on the RNA. By default, nodes are colored according to nucleotide identity.
@@ 5 Conclusion
In this chapter we demonstrated the use of rnaglib for convenient access to annotated RNA 3D structures and its use in a property prediction deep learning pipeline. We hope this illustration will lower the barrier to entry into the emerging field of geometric deep learning on RNA 3D structures and help inform novel RNA designs and uncover new structure-function relationships from 3D structure datasets. Finally, we point the reader again to the package documentation [^4] https://rnaglib.readthedocs.io/en/latest/ for a complete and up to date description of functionalities, including modules for self-supervised training and evaluation, and offer an invitation to contribute new features to the tool.
[…]